The stator of a three phase induction motor operates from a supply whose frequency is 50 Hz. Calculate the number of stator poles and total number of slots 1n the stator.


1
Expert's answer
2021-08-30T01:55:12-0400

The number of stator poles is


"p=\frac{2·60f}{n}=\frac{6000}{n},"

where n - the speed of the motor, the number of rotations per minute, or RPM.


The total number of slots in the stator:


"s=3mp=3·3·\frac{6000}{n}=\frac{54000}{n}."